If a, b, c, d,e are five consecutive positive even integers. There average is?
If a, b, c, d, e, are five consecutive positive odd integers, There average is?
Please let me know this answer urgently.

Answer by maheswari2188(4)

The numbers are a, b, c, d, e, f
As they are consecutive even integers they can be expressed as a, a+2, a+4, a+6, a+8.
Average = sum of the integers / Total Number of integers
ie, A = (a + a+2 + a+4 + a+6 + a+8)/ 5
A = (5a + 20)/ 5
A = 5(a + 4)/ 5 = a+4
2nd Part : same as above, as the difference between two consecutive odd integers are also 2.